Use the pythagorean Theorem to find the missing length when b=16 and c=20
Answer by checkley77(12844)

A^2+B^2=C^2
A^2+16^2=20^2
A^2+256=400
A^2=400-256
A^2=144
A=SQRT144
A=12 ANS.
PROOF:
12^2+16^2=20^2
144+256=400
400=400